Determine whether the variation between the indicated quantities is direct or inverse then explain why.
1) the radius of a circle and it’s area
2) the speed of a passenger train and the time required to travel between two cities on that train
3) if m varies directly as p, and m=36 when p=9, find m when p is 5

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

The larger the radius the larger the area, direct
the faster the speed the shorter the time, inverse
m=kp so 36=k*9 and k=4
so when p=5 m=4*5=20

Answer by ikleyn(52835)

.

Correction to the post by @Boreal :

    Variation between the radius of a circle and its area in NEITHER direct NOR inverse.


    It is the OTHER type of variation.